OpenStudy (anonymous):

Simplify 4 fourth root of 80

jimthompson5910 (jim_thompson5910):

What is 1^4 equal to

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

no, 1^4 = 1*1*1*1 = 1

jimthompson5910 (jim_thompson5910):

similarly, 2^4 = 2*2*2*2 = 16

jimthompson5910 (jim_thompson5910):

3^4 = 3*3*3*3 = 81 since we overshoot 80, we stop here so we only need to see if 2^4 = 16 goes into 80 1^4 = 1 goes into 80, but we ignore it since it's a trivial factor

jimthompson5910 (jim_thompson5910):

So does 16 go into 80?

jimthompson5910 (jim_thompson5910):

In other words, is 80/16 a whole number?

OpenStudy (anonymous):

ummm no?

jimthompson5910 (jim_thompson5910):

actually, 80/16 is a whole number since 80/16 = 5

jimthompson5910 (jim_thompson5910):

so this tells us that 80 = 16*5

OpenStudy (anonymous):

 Your 1st question

jimthompson5910 (jim_thompson5910):

why is this important? This is important because we factored 80 into two factors where one factor is a perfect 4th power This then allows us to say this \[\Large \sqrt[4]{80}\] \[\Large \sqrt[4]{16*5}\] \[\Large \sqrt[4]{16}*\sqrt[4]{5}\] \[\Large \sqrt[4]{2^4}*\sqrt[4]{5}\] \[\Large 2\sqrt[4]{5}\]

jimthompson5910 (jim_thompson5910):

So \[\Large \sqrt[4]{80} = 2\sqrt[4]{5}\]

OpenStudy (anonymous):

\[4 \sqrt[4]{80}\] this is the question

jimthompson5910 (jim_thompson5910):

so there's a 4 outside the 4th root?

jimthompson5910 (jim_thompson5910):

If that's the case, then multiply that outer 2 by 4 to get 8